The half life period of a radioactive element is 27.96 days .Calculate the time taken by a given sample to reduce to 1/8th of its initial activity.

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
half life period of a radioactive element is 27.96 days
t(1/2)=.693/k
k=.693/t(1/2)
k=2.303/T log (initial activity/final activity)
.693/27.96 =2.303/T *log(1/1/8)
T=84 days=the time taken by a given sample to reduce to 1/8th of its initial activity.